# Stability for the 2-D plane Poiseuille flow in finite channel

### 31 Dec 2023 v1

@@ 2.2. The improved transition threshold with Navier-slip boundary condition
$\|\omega\|_{L^{\infty}L^{2}}^{2}+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$ (4.2)
$\qquad+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
@@
$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}$ (4.3)
$\qquad+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
$\qquad+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
Proof.
The standard energy arguments give that
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle$
$=$ $\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}\omega\rangle=:\mathrm{RHS}.$
By Lemma 4.2, we have
$|k|$ $\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\lesssim|\mathrm{Im}\ \mathrm{RHS}|,$
which gives that
$|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}\lesssim|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}+|\mathrm{Im}\ \mathrm{RHS}|\lesssim|\mathrm{Im}\ \mathrm{RHS}|.$
Therefore, we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\frac{\nu}{2}\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
$\qquad\qquad+\frac{|k|}{2}(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\leq$ $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$
$\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\right)\leq C|\mathrm{RHS}|,$
where we have used Lemma 4.1 for $\omega(\pm 1)=0$.
The RHS is bounded as
$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
$\leq$ $\frac{1}{4}\left(\nu|k|^{2}\|\omega\|_{L^{2}}^{2}+\nu\|\partial_{y}\omega\|_{L^{2}}^{2}\right)+C\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2},$
or
$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
$\leq$ $\epsilon_{1}(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+C(\nu|k|^{2})^{-1/3}|k|^{2}\|f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{4}\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+C\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}.$
Taking $\epsilon_{1}\ll\min\{c,1\}$, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
$\qquad\qquad\qquad+|k|(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}}^{2}.$
Therefore, we get that
$\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
$\qquad\qquad\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}.$
Moreover, we have
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}$
$+|k|(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}}^{2}$
$+c_{0}(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}.$
Taking $c_{0}\ll 1$, we have
$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
$+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
The proof is completed. ∎
5. Nonlinear stability
In this section, we will give the proof of the main results, i.e., Theorem 1.1 and Theorem 1.2.
5.1. Nonlinear stability with non-slip boundary condition: Proof of Theorem 1.1
Recall that the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition read as
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \ u_{k}=(\partial_{y},-ik)\varphi_{k},\\ \varphi_{k}(t,k,\pm 1)=\partial_{y}\varphi_{k}(t,k,\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.1)
where
Let us recall that
$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$
Now we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
The global well-posedness of 2-D Navier-Stokes equation around the Poiseuille flow can be obtained by classical theory for Navier-Stokes equation. Here we only focus on the stability estimates. It should be noted that the case of $\nu|k|^{2}\geq 2$ is trivial by using (3.3) and we only consider the case of $\nu|k|^{2}\leq 2$.
Let us first consider the zero modes. If $k=0$, by the divergence-free condition, we have $\partial_{y}(u_{2})_{0}=0$, which implies by $(u_{2})_{0}|_{y=\pm 1}=0$ that $(u_{2})_{0}=0$. Therefore, the equation for zero modes is reduced to
By the similar arguments as for 2-D Couette flow [10], we have
$\mathcal{F}_{0}^{2}\lesssim\nu^{-1}\|(f_{2})_{0}\|_{L^{2}L^{2}}^{2}+\|\overline{\omega_{0}}\|_{L^{2}}^{2}.$ (5.2)
For $\mathcal{F}_{k}$, by Proposition 3.1, one has
$\mathcal{F}_{k}\lesssim|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}.$
For $k\not=0$, it holds that
$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$ (5.3)
$\leq$ $\sum_{l\in\mathbb{Z}}\left\|(u_{2})_{l}\right\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{\infty}L^{2}})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|^{\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\nu^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
and
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.4)
$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}.$
It remains to estimate the terms on the right-hand side of (5.9).
For $y\in[-1,0]$, it holds that $(1-|y|)=(1+y)\leq(1+y)(1-y)=1-y^{2}$; and for $y\in[0,1]$, $(1-|y|)=(1-y)\leq(1+y)(1-y)=1-y^{2}$, thus $1-y^{2}\geq 1-|y|$. For $l\not=0,k$, it holds that $|k|\lesssim|l||k-l|$, then
$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\left\|\frac{(u_{1})_{l}}{(1-y^{2})^{\frac{1}{2}}}\right\|_{L^{2}L^{\infty}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\partial_{y}(u_{1})_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\omega_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}(\nu|l|)^{-\frac{1}{4}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
and
$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\mathcal{F}_{0}\mathcal{F}_{k}.$
Thus, we have
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$ (5.5)
Combining (5.2), (5.3) and (5.5), we have
$\mathcal{F}_{k}\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
$\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l},$
$\mathcal{F}_{0}\lesssim$ $\|\overline{\omega}_{0}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}\setminus\{0\}}\mathcal{F}_{l}\mathcal{F}_{-l},$
which gives that
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\lesssim\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}.$
Due to $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$, then it holds that
$\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
If $\epsilon_{0}$ is suitably small, one can deduce from continuous argument that
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
The proof is completed. ∎
5.2. Nonlinear stability with Navier-slip boundary condition: Proof of Theorem 1.2
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ik(f_{1})_{k}-\partial_{y}(f_{2})_{k},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \varphi_{k}(\pm 1)=\omega_{k}(\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.6)
where
@@ 5. Nonlinear stability
$\mathcal{E}_{k}=(\nu|k|^{2})^{\frac{1}{6}}\|\omega_{k}\|_{L^{2}L^{2}}+\|\omega_{k}\|_{L^{\infty}L^{2}}+|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}},$
@@
$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}},$
where we have used the fact that $\min\{(\nu^{-1}|k|^{4})^{\frac{1}{3}},\nu^{-1}\}\leq(\nu|k|^{-1})^{-\frac{2}{3}}$.
@@
$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}\leq\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{\infty}}\|\omega_{k-l}\|_{L^{\infty}L^{2}}$ (5.8)
$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\mathcal{E}_{k-l}$
$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{2}L^{2}})^{\frac{1}{2}}\mathcal{E}_{k-l}$
@@
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.9)
$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}.$
@@
$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{2}}\mathcal{E}_{l}\nu^{-\frac{1}{6}}|k-l|^{-\frac{1}{3}}\mathcal{E}_{k-l}$
$\lesssim$ $\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\mathcal{E}_{l}\mathcal{E}_{k-l}$

### 03 Mar 2024 v2

@@ 4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow
$\left\{\begin{array}[]{lll}(\partial_{t}+\mathcal{L}_{k})\omega_{H}^{(2)}=\nu\partial_{y}^{2}\omega_{H}^{(1)}+4\nu|k|^{2}y^{2}t^{2}\omega_{H}^{(1)}+|\nu k|^{1/2}\omega_{H}^{(1)}\\ \qquad\qquad\qquad\quad+2ik(g(t,y)\varphi_{H}^{(0)}-\varphi_{H}^{(1)}),\\ (\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(2)}=\omega_{H}^{(2)},\ \omega_{H}^{(2)}|_{t=0}=0,\ \omega_{H}^{(2)}|_{y=\pm 1}=0,\end{array}\right.$
Following the arguments as in Lemma 4.2 of [9], we have
$|k|\|(\partial_{y},|k|)\varphi_{H}^{(1)}\|_{L^{2}L^{2}}^{2}\lesssim\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}.$ (4.8)
Next, we estimate $\|(\partial_{y},|k|)\varphi_{H}^{(2)}\|_{L^{2}L^{2}}$. Let us recall from Lemma 4.2 in [9] (here we set $k_{1}=k$) that
$\partial_{y}^{2}\omega^{(1)}_{H}$ $=\partial_{y}\omega_{1}^{(1)}+2ikt\omega_{H}^{(1)}+2ikty(\omega_{1}^{(1)}+2ikty\omega^{(1)}_{H}),$
$\partial_{y}^{2}\omega_{H}^{(1)}$ $+4|k|^{2}y^{2}t^{2}\omega^{(1)}_{H}=\partial_{y}\omega_{1}^{(1)}+2ikt\omega^{(1)}_{H}+2ikty\omega_{1}^{(1)},$
$\omega_{1}=X{\omega}^{(0)}_{H},\ \omega_{1}^{(1)}=X\omega^{(1)}_{H},\ X=\partial_{y}-2ikyt.$
$\phi_{1}^{(1)}=g(t,y)\varphi_{H}^{(0)}-\varphi_{H}^{(1)},$
and we have
$\partial_{t}\omega^{(2)}_{H}-\nu(\partial_{y}^{2}-|k|^{2})\omega_{H}^{(2)}+ik\big{(}(1-y^{2})\omega_{H}^{(2)}+2\varphi^{(2)}_{H}\big{)}$
$=\nu(\partial_{y}\omega_{1}^{(1)}+2ikt\omega^{(1)}_{H}+2ikty\omega_{1}^{(1)})+|\nu k|^{1/2}\omega_{H}^{(1)}+2ik\phi_{1}^{(1)}.$
Applying Lemma 4.1 with $f_{2}=-\nu\omega_{1}^{(1)}$, $f_{4}=-2\nu ikty\omega_{1}^{(1)}-2\nu ikt\omega_{H}^{(1)}-(\nu|k|)^{\frac{1}{2}}\omega_{H}^{(1)}$ and $f_{3}=-2ik\phi_{1}^{(1)}$ that
$\|(\partial_{y},|k|)\varphi_{H}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim$ $\nu^{-1}|k|^{-1}\|\nu\omega_{1}^{(1)}\|_{L^{2}L^{2}}^{2}+|k|^{-2}\|k(\partial_{y},|k|)\phi_{1}^{(1)}\|_{L^{2}L^{2}}^{2}$
$+\nu^{-\frac{1}{2}}|k|^{-\frac{3}{2}}\bigg{(}\|\nu kty\omega_{1}^{(1)}\|_{L^{2}L^{2}}^{2}+\|(|\nu kt|+(\nu|k|)^{\frac{1}{2}})\omega_{H}^{(1)}\|_{L^{2}L^{2}}^{2}\bigg{)}$
$\lesssim$ $|k|^{-1}\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
Therefore, we conclude that
$|k|\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2},$
which completes the proof for case of $\nu|k|^{2}\leq 2$.
If $\nu|k|^{2}\geq 2$, then it holds that
$|k|^{2}\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\nu|k|^{2}\|\omega_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\omega_{0}\|_{L^{2}}^{2},$
which completes that proof. ∎
5. Nonlinear stability
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}+\mathcal{L}_{k}\omega_{k}=-ik(f_{1})_{k}-\partial_{y}(f_{2})_{k},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\\ \varphi_{k}(\pm 1)=\omega_{k}(\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.1)
where
$\mathcal{L}_{k}:=-\nu(\partial_{y}^{2}-k^{2})+ik(1-y^{2})+2ik(\partial_{y}^{2}-|k|^{2})^{-1}$
@@ 5. Nonlinear stability
$\mathcal{E}_{k}=$ $(\nu|k|)^{\frac{1}{4}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$
$+\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+|k|^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}},$
@@
$\mathcal{E}_{k}\lesssim\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}.$
where we have used the fact that $|k|\min\{(\nu|k|)^{-\frac{1}{4}},(\nu|k|^{2})^{-\frac{1}{2}}\}\leq\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}$.
@@
$\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$ (5.3)
$\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\|_{L^{2}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\partial_{y}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\mathcal{E}_{k-l}$
$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{2}L^{2}})^{\frac{1}{2}}\mathcal{E}_{k-l}$
@@
$\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.4)
$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}.$
@@
$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\partial_{y}(u_{1})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}$
×$\times\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}^{\frac{5}{6}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}^{\frac{1}{6}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{2}}\mathcal{E}_{l}(\nu|k-l|)^{-\frac{1}{4}\cdot\frac{5}{6}}(\nu^{\frac{1}{2}}|k-l|)^{-\frac{1}{6}}\mathcal{E}_{k-l}$
$\lesssim$ $\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{8}}\mathcal{E}_{l}\mathcal{E}_{k-l}$